There is a 40% chance that it will be cloudy tomorrow.
If it is cloudy, there is a 79% chance that it will rain.
What is the probability that it will rain?
Answer by ikleyn(53751) 
 .
There is a 40% chance that it will be cloudy tomorrow.
If it is cloudy, there is a 79% chance that it will rain.
What is the probability that it will rain?
~~~~~~~~~~~~~~~~
To make the problem logically complete and ready for a mathematical solution,
I will add an additional assumption, which is consistent with the common sense and with Physics laws.
This addition is
If it is not cloudy, then there is no rain.
Now I am ready to solve the problem, following the rules of probability.
The formula for the probability is
P(rain) = P(cloudy)*0.79 + P(not cloudy)*0 = I continue = 0.40*0.79 + (1-0.40)*0 = 0.316.
ANSWER. The probability is 0.316 that it will rain.
Solved.
|
|
|